# Dance Like a Chicken: Low-Rank Stylization for Human Motion Diffusion

### v1

@@ Abstract
Figure 1: LoRA-MDM generates stylized motions using Low-Rank Adaptation of the Motion Diffusion Model (MDM). Left: motions generated using the base model. Middle: stylized generation using LoRA-MDM. Left: Style mixing.
@@ 1 Introduction
Equipped with this insight, a simple solution, of employing Low-Rank Adaptation (LoRA) [26], becomes clear; LoRA augments the weights of the prior in a low-ranked manner, which can be intuitively thought of as low frequency deformations to the target distribution manifold. Using LoRA, LoRA-MDM learns to embed the style within the output domain of the model, which can then be used to synthesize stylized motions in the context of different textual prompts. As we demonstrate, this offers two advantages: First, low-ranked augmentation diminishes forgetting, which in turn increases generation quality for stylized actions not seen during training. For example, directly fine-tuning a generated walking motion to walk like a chicken has little (or negative) influence on how the prior will generate kicks. Instead, our low-ranked tuning and structure preservation loss encourage the model to also kick like a chicken, since the changes affect the entire manifold. Second, because the generated stylized motions are in-distribution for the model, advanced operations such as stylized motion editing and style blending are supported, in sharp contrast to traditional motion style-transfer method.
@@ 3.1 Preliminaries
We follow the common practice in motion diffusion genration and using the Denoising Diffusion Probabilistic Models (DDPM) formulation [24]. The diffusion process consists of forward and backward processes. During training, in the forward pass, we add noise $\epsilon$ to the clean data $x_{0}=x$ according to:
@@
During inference, we iteratively apply $G(x_{t},t,\mathbf{c})$ from $t=T$ down to $t=0$ to recover $\hat{\mathbf{x}}_{0}$.
@@ 4 Experiments
Figure 3: Qualitative Results. The LoRA adapter is trained on a handful of style reference motions (blue, left). Generated motions (orange, right) adhere to both the text prompt and style. Note how the prior fits Capoeira moves to ”fighting” and frog leaps to ”jumping” to better fit both the text prompt and the learned style.
@@ 4.1 Evaluation Setting
Real 90.9% 97.7% 97.7% 79.6% 0.49 2.96 0.057 9.05
LoRA-MDM (ours) 27.1% 42.5% 39.8% 80.0% 0.32 3.01 0.055 11.07
@@ 6 Discussion
Our experimentation also highlights some of the well known problems of style transfer in general, and in the motion domain specifically. Similar to the image domain, the metrics used are unreliable. There is no metric beside the subjective one to measure the quality of an airplane style kicking motion. All automatic evaluators are either sensitive to the action performed and are unfamiliar with most styles, or vice versa. In both cases, stylized actions are out-of-domain and poorly evaluated. This problem is even more acute for the motion domain, as data is more scarce. There are only a handful of style datasets available, consisting few actions, and demonstrating unclear styles. For example, the largest dataset, which we used, includes styles that are actually actions (e.g., punching) or are hard to define or distinguish (e.g., Star and Monk). Indeed, most literature employs only a subset of the styles for training and evaluation since most styles in the dataset are confusing (see Appendix C.2) In the future, we hope this work helps set common practices for applying LoRA, and facilitate applying it to other motion domains, extending motion generation capabilities in other underrepresented axes except style.
@@ B.1 Motion Representation
The 100STYLE dataset  [39] consists of motion capture clips of seven movement types (Backwards Running, Backwards Walking, Forwards Running, Forwards Walking, Idling, Sidestep Running, and Sidestep Walking) performed in 100 different styles. The dataset was post-processed by Zhong et al. [73], who primarily retargeted to the SMPL  [37] skeleton and segmented into clips. Some of the released styles were poorly processed and were excluded from all our experiments. See Table 3 for details.

### v2

@@ Abstract
Figure 1: LoRA-MDM generates stylized motions using Low-Rank Adaptation of the Motion Diffusion Model (MDM). Left: motions generated using the base model. Middle: stylized generation using LoRA-MDM. Right: Style mixing.
@@ 1 Introduction
Equipped with this insight, a simple solution, of employing Low-Rank Adaptation (LoRA) [26], becomes clear; LoRA augments the weights of the prior in a low-ranked manner, which can be intuitively thought of as low frequency deformations to the target distribution manifold. Using LoRA, LoRA-MDM learns to embed a style within the output domain of the model, which can then be used to synthesize stylized motions in the context of different textual prompts. As we demonstrate, this offers two advantages: First, low-ranked augmentation diminishes forgetting, which in turn increases generation quality for stylized actions not seen during training. For example, directly fine-tuning a generated walking motion to walk like a chicken has little (or negative) influence on how the prior will generate kicks. Instead, our low-ranked tuning and structure preservation loss encourage the model to also kick like a chicken, since the changes affect the entire manifold. Second, because the generated stylized motions are in-distribution for the model, advanced operations such as stylized motion editing and style blending are supported, in sharp contrast to traditional motion style-transfer method.
@@ 3.1 Preliminaries
We follow the common practice in motion diffusion generation and use the Denoising Diffusion Probabilistic Models (DDPM) formulation [24]. The diffusion process consists of forward and backward processes. During training, in the forward pass, we add noise $\epsilon$ to the clean data $x_{0}=x$ according to:
@@
During inference, $G(x_{t},t,\mathbf{c})$ is iteratively applied from $t=T$ down to $t=0$ to recover $\hat{\mathbf{x}}_{0}$.
@@ 4 Experiments
Figure 3: Qualitative Results. The LoRA adapter is trained on a handful of style reference motions (blue, left). Generated motions (orange, right) adhere to both the text prompt and style. Notably, the dynamic style “drunk” applied to “raise hands” reveals semantic style transfer, and the “chicken” style with “dance wildly” yields smooth and coherent motion, demonstrating non-trivial generalization. Also, the prior fits Capoeira moves to ”fighting” and frog leaps to ”jumping” to better fit both the text prompt and the learned style.
@@ 4.1 Evaluation Setting
Real 90.9% 97.7% 97.7% 79.6% 2.96 0.057 9.05
LoRA-MDM (ours) 27.3% 43.0% 70.3% 79.3% 0.31 3.03 0.054 8.31
@@ 6 Discussion
Our experimentation also highlights some of the well known problems of style transfer in general, and in the motion domain specifically. Similar to the image domain, the metrics used are unreliable. There is no metric beside the subjective one to measure the quality of an airplane style kicking motion. All automatic evaluators are either sensitive to the action performed and are unfamiliar with most styles, or vice versa. In both cases, stylized actions are out-of-domain and poorly evaluated. This problem is even more acute for the motion domain, as data is more scarce. There are only a handful of style datasets available, consisting few actions, and demonstrating unclear styles. For example, the largest dataset, which we used, includes styles that are actually actions (e.g., punching) or are hard to define or distinguish (e.g., Star and Monk). Indeed, most literature employs only a subset of the styles for training and evaluation since most styles in the dataset are confusing (see Appendix C.2). In the future, we hope this work helps set common practices for applying LoRA, and facilitate applying it to other motion domains, extending motion generation capabilities in other underrepresented axes except style.
@@ B.1 Motion Representation
The 100STYLE dataset  [39] consists of motion capture clips of seven movement types (Backwards Running, Backwards Walking, Forwards Running, Forwards Walking, Idling, Sidestep Running, and Sidestep Walking) performed in 100 different styles. The dataset was post-processed by Zhong et al. [73], who primarily retargeted to the SMPL  [37] skeleton and segmented into short clips, each contains 196 frames. Some of the released styles were poorly processed and were excluded from all our experiments. See Table 3 for details.